you are evaluating purchasing the rights to a project that will generate after tax expected cash flows of $84k at the end of each of yhe next five years, plus an additional $1,000k at the end of the fifth year as the final cash flow. you can purchase this project for $608 k. if your firms cost of capital (aka required rate of return) is 12.5% what is the NPV of this project?
Answer #1
PV of annual after tax cash inflows = 84000*(1.125^5-1)/(0.125*1.125^5) = $   2,99,087.74
PV of addl. cash inflow at EOY 5 = 1000000/1.125^5 = $   5,54,928.96
Total PV of cash inflows $   8,54,016.70
Less: Initial cash outflow $   6,08,000.00
NPV $   2,46,016.70